OpenStudy (anonymous):

2) 2/5x^2 -x+ 19/15 =0

OpenStudy (anonymous):

lol

OpenStudy (just_one_last_goodbye):

anyways lets see how i can help

OpenStudy (just_one_last_goodbye):

is the first 2 part of the equation or no?

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }x^2-x+\frac{ 19 }{ 15 }=0\]

OpenStudy (anonymous):

I hope that makes it better

OpenStudy (just_one_last_goodbye):

Much better ^-^

OpenStudy (anonymous):

So I know A= 2/5 B= -1 C= 19/15

OpenStudy (just_one_last_goodbye):

Decimal answer correct?

OpenStudy (anonymous):

yes

OpenStudy (just_one_last_goodbye):

and Solving for x?

OpenStudy (just_one_last_goodbye):

x = 1.455485

OpenStudy (anonymous):

it says "Use the quadratic formula to solve the equation." "A. The solution set is ____" (simplify your answer. type an exact answer, ysing radicals as needed. Use intefers or fractions for any numbers in the expression. Use a comma to separate answers as needed.) "B. the solution is (emptyset)"

OpenStudy (just_one_last_goodbye):

~eats pizza while waiting~

OpenStudy (anonymous):

lol my b

OpenStudy (just_one_last_goodbye):

~passes skittles~

OpenStudy (anonymous):

omg. skittles with coffee

OpenStudy (anonymous):

yes

OpenStudy (just_one_last_goodbye):

^

OpenStudy (just_one_last_goodbye):

Oh quadratic thought it was a normal equation sorry

OpenStudy (anonymous):

lol, the only unit i have left is quadratic and I haven't done it in 2 years

OpenStudy (just_one_last_goodbye):

hmm idk it's not giving me an answer

OpenStudy (anonymous):

okay it's okay, let's move on to the next question?

OpenStudy (just_one_last_goodbye):

sure

OpenStudy (anonymous):

.6x^2 = 1.2x+0.6

OpenStudy (anonymous):

\[0.6x^2-1.2x-0.6=0\]

OpenStudy (just_one_last_goodbye):

wait is it an addition problem or subtraction?

OpenStudy (anonymous):

I just flipped it so it's in standard form. it's still quadratic. just now it's all on one side.

OpenStudy (anonymous):

...

OpenStudy (just_one_last_goodbye):

Open this attachment up, It should help ^-^

OpenStudy (anonymous):

what site did you use?

OpenStudy (misty1212):

HI!!

OpenStudy (just_one_last_goodbye):

. - . . _ .

OpenStudy (misty1212):

\[\frac{ 2 }{ 5 }x^2-x+\frac{ 19 }{ 15 }=0\] is annoying multiply by 15 and get \[6x^2-15x+19=0\] then it will be much easier

OpenStudy (misty1212):

you can use the quadratic formula or if you are a factoring wiz you can try that

OpenStudy (anonymous):

yeah. I'm on here, asking for help. I'm not a factoring wiz.

OpenStudy (misty1212):

ok well is it really \[\frac{ 2 }{ 5 }x^2-x+\frac{ 19 }{ 15 }=0\]?

OpenStudy (misty1212):

because that equation has no real solutions, maybe there is a typo there

OpenStudy (anonymous):

yes. that's how it's written on my problem

OpenStudy (anonymous):

well one of the answers is there is no solution

OpenStudy (misty1212):

then either you are supposed to come up with two ugly complex solutions, or say "no solution"

OpenStudy (anonymous):

so you sorta answered my question

OpenStudy (misty1212):

ok fine in any case you really want to multiply by 15 first and start with \[6x^2-15x+19=0\] then, when you use the quadratic formula, you will see that there is no answer

OpenStudy (misty1212):

you do not want to use decimals you do not want to use fractions you want to work with integers only

OpenStudy (anonymous):

okay, thank you